OpenStudy (anonymous):

etermine whether the graph of y = x2 + 4x − 7 has a maximum or minimum point, then find the maximum or minimum value. A. Minimum; −2 B. Maximum; −2 C. Maximum; −11 D. Minimum; −11

OpenStudy (anonymous):

Minimum = -2

OpenStudy (savvy):

minimum: -11

OpenStudy (anonymous):

x = -2, y = -11

OpenStudy (savvy):

it lies at x=-2

OpenStudy (anonymous):

you can complete the square pretty easily here x^2+4x-7= x^2+4x+4-4-7= (x+2)^2-11 a(x-h)^2+k where (h,k) is the vertex so your vertex is at (-2, -11) -11 is the min value because a>0